# YOLO-TLA: An Efficient and Lightweight Small Object Detection Model based on YOLOv5

### v1

@@ IV-B Evaluation metrics
[…]
@@
IV-C Evaluations on tiny object detection layer
[…]
TABLE II: Detection results of tiny object detection layers.
[…]
YOLOv5s 67.7% 50.3% 55.7% 36.1% 0.577 16.6 7.23
@@ IV-C Evaluations on tiny object detection layer
To highlight the efficacy of the proposed tiny object detection layer, we utilize result images and head network heat maps from both the YOLOv5s and YOLOv5s-Tiny models. Fig.7 shows (a) the original image, (b) inference results of YOLOv5s, and (c) inference results of YOLOv5s-Tiny, indicating a superior ability of YOLOv5s-Tiny in detecting small objects. Fig.8 presents three feature heat maps from the YOLOv5s neck network, corresponding to the detection layers for large, medium, and small objects, while Fig.9 illustrates the corresponding layers from YOLOv5s-Tiny, including the tiny object detection layer. Notably, Figs.9 and 9 reveal that the tiny object detection layer in YOLOv5s-Tiny offers more detailed attention to small objects.
@@
Figure 8: Heat map of YOLOv5s neck network outputs.
Figure 9: Heat map of YOLOv5s-Tiny neck network outputs.
IV-D Evaluations on lightweight convolution modules
In this study, four methods to lighten the YOLOv5 model are proposed. The first method, termed YOLOv5s-G1, replaces the C3 module in the backbone with a C3Ghost module. The second, YOLOv5s-G2, switches out all C3 modules in the network for C3Ghost modules. The third method, YOLOv5s-CC1, swaps the C3 module in the backbone for a C3CrossCovn module. To avoid confusion, the fourth method, which replaces all network C3 modules with C3CrossCovn modules, is named YOLOv5s-CC2. The experimental results for these four models are compiled in Table III.
TABLE III: Detection results of four lightweight convolution modules.
[…]
YOLOv5s 67.7% 50.3% 55.7% 36.1% 0.577 16.6 7.23
@@ IV-D Evaluations on lightweight convolution modules
The four lightweight improved models, particularly YOLOv5s-G1 and YOLOv5s-G2, have achieved reductions in parameters and computational demands compared to the baseline YOLOv5 model. However, YOLOv5s-G1 and YOLOv5s-G2 also experienced a notable decrease in detection accuracy. This drop is attributed to two issues: the loss of channel features due to the feature map splicing and the varying receptive fields of GhostCovn, leading to diminished representation capability. On a positive note, the YOLOv5s-CC1 model not only reduced its parameter count by 0.2M and computational complexity by 0.4GFLOPs but also improved accuracy, with a 1.9% increase in mAP@0.50 and a 1.6% increase in mAP@0.5:0.95. These results indicate the efficacy of the C3CrossCovn module in enhancing backbone network feature extraction, although it presents challenges for multi-scale feature fusion in the neck network and impacts accuracy due to the overall reduction in model parameters and complexity.
IV-E Evaluations on different attention mechanisms
YOLO-TLA employs GAM within its backbone network and compares it with other prevalent attention mechanisms such as CA, CBAM, and ECA in terms of their effects on the detection performance. As detailed in Table IV, each attention mechanism enhances various detection metrics of YOLOv5s to different extents. The YOLOv5s model with SKNet demonstrates a significant increase of 4.5% in mAP@0.50 and 4% in mAP@0.5:0.95, albeit at the cost of a nearly fivefold increase in model parameters and complexity, conflicting with the lightweight objective of this study. Conversely, the YOLOv5s model integrated with GAM achieves improvements of 3.5% in mAP@0.50 and 3.3% in mAP@0.5:0.95, with minimal impact on model size and complexity. Therefore, among the tested attention mechanisms, GAM stands out as the most effective in enhancing detection accuracy without substantially increasing model size, maintaining an optimal balance. The comparative impacts of these attention mechanisms on YOLOv5s are visually presented in Fig.10, using metrics like mAP@0.50, mAP@0.5:0.95, parameter count, and computational complexity.
TABLE IV: Detection results of seven attention mechanisms.
[…]
YOLOv5s 67.7% 50.3% 55.7% 36.1% 0.577 16.6 7.23
@@ IV-E Evaluations on different attention mechanisms
IV-F Discussions
To illustrate the effects of each enhancement on model performance, the most effective improvement method from each category is chosen in this section. Specifically, YOLOv5s-TL adds both a tiny object layer and a lightweight improvement, based on the YOLOv5s-CC1 model. YOLOv5s-TA combines a tiny object detection layer with GAM but retains the standard C3 module. Finally, YOLO-TLAs, the model proposed in this paper, integrates the tiny object detection layer, the lightweight C3CrossCovn module, and GAM. The outcomes of these various modifications are detailed in Table V.
TABLE V: Detection results of different improved models.
[…]
YOLOv5s 67.7% 50.3% 55.7% 36.1% 0.577 16.6 7.23
@@ IV-F Discussions
YOLO-TLA 71.2% 57.3% 60.3% 40.1% 0.635 25.3 9.49
Compared to YOLOv5s-Tiny, YOLOv5s-TL shows a 0.6% improvement in mAP@0.50 and a 0.5% improvement in mAP@0.5:0.95, along with a reduction of 0.19M in parameters. YOLOv5s-TA demonstrates a 2.1% higher mAP@0.50 and a 2.2% higher mAP@0.5:0.95 compared to YOLOv5s-Tiny. Relative to YOLOv5s-TA, YOLO-TLAs reduces parameters by 0.21M and computational demands by 0.2GFLOPs, while achieving a 0.7% improvement in mAP@0.50 and a 0.4% improvement in mAP@0.5:0.95. These results reveal that the lightweighting approach not only lowers model complexity but also enhances detection performance. While improvements in accuracy often lead to increased computational and parameter requirements, the lightweighting strategy implemented here effectively limits these increases. The experimental results validate that each proposed improvement positively contributes to the enhanced performance of the YOLOv5s model and confirm that these improvements are not mutually exclusive but rather complementary.
YOLO-TLAs demonstrates notable performance enhancements over YOLOv5s, with increases of 4.6% in mAP@0.50 and 4% in mAP@0.5:0.95, and a parameter increase of just 2.26M. Applying the same improvement method to the larger YOLOv5m model resulted in the YOLO-TLAm model. This model shows significant gains, with a 1.7% increase in mAP@0.50 and a 1.9% increase in mAP@0.5:0.95, compared to YOLOv5m, but at the cost of an additional 6.36M parameters. YOLOv5 adjusts model size by modifying network dimensions, particularly by varying the repetition of modules. Larger models like YOLO-TLAm generally see more substantial parameter and computational increases compared to similar-sized YOLOv5 models, without corresponding large gains in accuracy. Table VI presents a performance comparison of different YOLO-TLA models with their YOLOv5 equivalents.
TABLE VI: Detection results of different model sizes.
[…]
YOLOv5s 67.7% 50.3% 55.7% 36.1% 0.577 16.6 7.23
YOLOv-TLAs 71.2% 57.3% 60.3% 40.1% 0.635 25.3 9.49
YOLOv5m 70.7% 58.2% 63.1% 43.4% 0.638 48.9 21.17
YOLOv-TLAm 74.6% 60.3% 64.8% 45.3% 0.667 73.1 27.53
@@
IV-G Comparison with state-of-the-arts
Table VII reports the detailed comparison of YOLO-TLAs with other state-of-the-art detection methods. YOLO-TLAs exhibits a 0.3% improvement in mAP@0.5:0.95 over YOLOx-s, alongside a reduction of 1.5GFLOPs in computational demand, but with a parameter increase of 0.49M. YOLO-TLAm also outperforms its counterpart YOLOx-m, showing a 0.4% better mAP@0.5:0.95 and a 0.7GFLOPs decrease in computation. Moreover, YOLO-TLAs significantly betters EfficientDet-D1 by 7.5% in mAP@0.5:0.95. While the network architecture of EfficientDet-D1 is lightweight, its simplicity results in a trade-off with accuracy.
TABLE VII: Comparative results of state-of-the-arts.
@@ IV-G Comparison with state-of-the-arts
YOLOv6n [11] 37.2% 11.4 4.70
YOLOv8n 36.9% 8.7 3.20
EfficientDet-D1 [23] 32.6% 6.1 6.60
YOLO-TLAs 40.1% 25.3 9.49
YOLO-TLAm 45.3% 73.1 27.53
V Conclusion
This study tackles prevalent challenges in object detection communities and introduces a novel method, YOLO-TLA. This method demonstrates superior detection performance compared to the baseline YOLOv5, particularly in accurately identifying small objects. YOLO-TLA enhances YOLOv5 by integrating a tiny object detection layer into its neck network, and by incorporating a global attention mechanism to increase accuracy. To balance enhanced detection with model efficiency, a lightweight strategy is employed, incorporating the C3CrossCovn module in the backbone network to decrease model complexity. This strategy not only lowers complexity but also boosts accuracy, as confirmed by experimental results. Furthermore, we also successfully applies these improvements to the larger YOLOv5m model, creating the YOLO-TLAm model, which outperforms YOLOv5m in accuracy and stability. The results validate that the proposed enhancements are effective for larger models as well.

### v2

@@ 4.2 Evaluation metrics
Given the extensive range of object categories in the MS COCO dataset employed in our experiment, it is impractical to display precision and recall for each category individually. As a result, we presents the metrics as averages across all categories, with average precision and recall represented by Precisionall and Recallall, respectively. The formulas are expressed as follows,
@@
4.3 Evaluations on tiny object detection layer
Table 2: Detection results of tiny object detection layers.
Models Precisionall Recallall mAP@0.50 mAP@0.5:0.95 F1 GFLOPs Parameters (M)
YOLOv5s [12] 67.7% 50.3% 55.7% 36.1% 0.577 16.6 7.23
@@ 4.3 Evaluations on tiny object detection layer
Figure 8: Output heatmaps of YOLOv5s neck network.
Figure 9: Output heatmaps of YOLOv5s-Tiny neck network.
By incorporating tiny object detection layers in the neck network, YOLOv5s-Tiny outperforms the baseline YOLOv5s in several metrics, as detailed in Table 2. This enhancement has led to a 1% gain in Precisionall, a 2.2% gain in Recallall, a 1.8% gain in mAP@0.50, and a 1.4% gain in mAP@0.5:0.95. However, this has resulted in higher computational and parameter requirements, likely due to the additional convolutional modules in the neck network that create larger feature maps.
To highlight the efficacy of the proposed tiny object detection layer, we utilize result images and head network heat maps from both the YOLOv5s and YOLOv5s-Tiny models. Fig.7 shows (a) the original image, (b) detection results of YOLOv5s, and (c) detection results of YOLOv5s-Tiny, indicating a superior ability of YOLOv5s-Tiny in detecting small objects. Fig.8 presents three feature heat maps from the YOLOv5s neck network, corresponding to the detection layers for large, medium, and small objects, while Fig.9 illustrates the corresponding layers from YOLOv5s-Tiny, including the tiny object detection layer. Notably, Figs.9 and 9 reveal that the tiny object detection layer in YOLOv5s-Tiny offers more detailed attention to small objects.
4.4 Evaluations on lightweight convolution modules
In this study, four methods to lighten the YOLOv5 model are proposed. The first method, termed YOLOv5s-G1, replaces the C3 module in the backbone with a C3Ghost module. The second method, YOLOv5s-G2, switches out all C3 modules in the network for C3Ghost modules. The third method, YOLOv5s-CC1, swaps the C3 module in the backbone for a C3CrossCovn module. To avoid confusion, the fourth method, which replaces all network C3 modules with C3CrossCovn modules, is named YOLOv5s-CC2. The experimental results for these four models are compiled in Table 3.
Table 3: Detection results of four lightweight convolution modules.
Models Precisionall Recallall mAP@0.50 mAP@0.5:0.95 F1 GFLOPs Parameters (M)
YOLOv5s [12] 67.7% 50.3% 55.7% 36.1% 0.577 16.6 7.23
@@ 4.4 Evaluations on lightweight convolution modules
Table 4: Detection results of seven attention mechanisms.
Models Precisionall Recallall mAP@0.50 mAP@0.5:0.95 F1 GFLOPs Parameters (M)
YOLOv5s [12] 67.7% 50.3% 55.7% 36.1% 0.577 16.6 7.23
@@
The four lightweight improved models, particularly YOLOv5s-G1 and YOLOv5s-G2, have achieved reductions in parameters and computational demands compared to the baseline YOLOv5 model. However, YOLOv5s-G1 and YOLOv5s-G2 also experienced a notable decrease in detection accuracy. Specifically, compared to the baseline model, YOLOv5s-G1 decreased the mAP@0.50 by 0.7%, mAP@0.5:0.95 by 0.6%, the number of parameters by 1.17M, and the computational complexity by 3.6GFLOPs. For the YOLOv5s-G2, its mAP@0.50 value decreased by 1.3% compared to YOLOv5s, mAP@0.5:0.95 decreased by 1.1%, the number of parameters decreased by 2.13M, and the computational complexity decreased by 5.5GFLOPs. These drops can be attributed to the loss of channel features due to the feature map splicing and the varying receptive fields of GhostCovn, leading to diminished representation capability. The C3CrossConv module includes CrossConv, which focuses more on the main regions of an image, often containing more detectable objects while selectively ignoring other regions with fewer objects. This feature extraction method somewhat reduces model complexity and computational cost without sacrificing accuracy. On a positive note, the YOLOv5s-CC1 model not only reduced its parameter count by 0.2M and computational complexity by 0.4GFLOPs but also improved accuracy, with a 1.9% increase in mAP@0.50 and a 1.6% increase in mAP@0.5:0.95. These results indicate the efficacy of the C3CrossCovn module in enhancing backbone network feature extraction, although it presents challenges for multi-scale feature fusion in the neck network and impacts accuracy due to the overall reduction in model parameters and complexity.
In summary, the application of C3CrossConv to the backbone network yielded gratifying results, achieving a dual effect: it not only reduces the number of parameters and computational cost of the model but also enhances the accuracy and efficiency. Therefore, we choose to apply C3CrossConv in the backbone network of YOLO-TLA to achieve optimization.
4.5 Evaluations on different attention mechanisms
YOLO-TLA employs GAM within its backbone network and compares it with other prevalent attention mechanisms such as CA, CBAM, and ECA in terms of their effects on detection performance. As detailed in Table 4, each attention mechanism enhances various detection metrics of YOLOv5s to different extents. The YOLOv5s model with SKNet demonstrates a significant increase of 4.5% in mAP@0.50 and 4% in mAP@0.5:0.95, albeit at the cost of a nearly fivefold increase in model parameters and complexity, conflicting with the lightweight objective of this study. Conversely, the YOLOv5s model integrated with GAM achieves improvements of 3.5% in mAP@0.50 and 3.3% in mAP@0.5:0.95, with minimal impact on model size and complexity. Therefore, among the tested attention mechanisms, GAM stands out as the most effective in enhancing detection accuracy without substantially increasing model size, maintaining an optimal balance. The comparative impacts of these attention mechanisms on YOLOv5s are visually presented in Fig.10, using metrics like mAP@0.50, mAP@0.5:0.95, parameter count, and computational complexity.
4.6 Discussions
To illustrate the effects of each enhancement on model performance, the most effective improvement method from each category is chosen in this section. Specifically, YOLOv5s-TL adds both a tiny object layer and a lightweight improvement, based on the YOLOv5s-CC1 model. YOLOv5s-TA combines a tiny object detection layer with GAM but retains the standard C3 module. Finally, YOLO-TLAs, the model proposed in this paper, integrates the tiny object detection layer, the lightweight C3CrossCovn module, and GAM. The outcomes of these various modifications are detailed in Table 5.
Table 5: Detection results of different improved models.
Models Precisionall Recallall mAP@0.50 mAP@0.5:0.95 F1 GFLOPs Parameters (M)
YOLOv5s [12] 67.7% 50.3% 55.7% 36.1% 0.577 16.6 7.23
@@ 4.6 Discussions
YOLO-TLA (Ours) 71.2% 57.3% 60.3% 40.1% 0.635 25.3 9.49
Table 6: Detection results of different model sizes.
Models Precisionall Recallall mAP@0.50 mAP@0.5:0.95 F1 GFLOPs Parameters (M)
YOLOv5s [12] 67.7% 50.3% 55.7% 36.1% 0.577 16.6 7.23
YOLOv-TLAs (Ours) 71.2% 57.3% 60.3% 40.1% 0.635 25.3 9.49
YOLOv5m [12] 70.7% 58.2% 63.1% 43.4% 0.638 48.9 21.17
YOLOv-TLAm (Ours) 74.6% 60.3% 64.8% 45.3% 0.667 73.1 27.53
@@
Table 7: Comparative results of YOLO-TLAs against other lightweight object detection models.
@@
YOLOv6n [11] 37.2% 11.4 4.70
YOLOv8n [17] 36.9% 8.7 3.20
EfficientDet-D1 [46] 32.6% 6.1 6.60
YOLO-TLAs (Ours) 40.1% 25.3 9.49
Table 8: Comparative results of YOLO-TLAm against other large-scale object detection models.
Models mAP@0.5:0.95 GFLOPs Parameters (M)
YOLOv5m [12] 43.4% 48.9 21.17
@@
YOLO-TLAm (Ours) 45.3% 73.1 27.53
Compared to YOLOv5s-Tiny, YOLOv5s-TL shows a 0.6% improvement in mAP@0.50 and a 0.5% improvement in mAP@0.5:0.95, along with a reduction of 0.19M in parameters. YOLOv5s-TA demonstrates a 2.1% higher mAP@0.50 and a 2.2% higher mAP@0.5:0.95 compared to YOLOv5s-Tiny. Relative to YOLOv5s-TA, YOLO-TLAs reduces parameters by 0.21M and computational demands by 0.2GFLOPs, while achieving a 0.7% improvement in mAP@0.50 and a 0.4% improvement in mAP@0.5:0.95. These results reveal that the lightweight approach not only lowers model complexity but also enhances detection performance. While improvements in accuracy often lead to increased computational and parameter requirements, the lightweight strategy implemented here effectively limits these increases. The experimental results validate that each proposed improvement positively contributes to the enhanced performance of the YOLOv5s model and confirm that these improvements are not mutually exclusive but rather complementary.
YOLO-TLAs demonstrates notable performance enhancements over YOLOv5s, with increases of 4.6% in mAP@0.50 and 4% in mAP@0.5:0.95, and a parameter increase of just 2.26M. Applying the same improvement method to the larger YOLOv5m model resulted in the YOLO-TLAm model. This model shows significant gains, with a 1.7% increase in mAP@0.50 and a 1.9% increase in mAP@0.5:0.95, compared to YOLOv5m, but at the cost of an additional 6.36M parameters. YOLOv5 adjusts model size by modifying network dimensions, particularly by varying the repetition of modules. Larger models like YOLO-TLAm generally see more substantial parameter and computational increases compared to similar-sized YOLOv5 models, without corresponding large gains in accuracy. Table 6 presents a performance comparison of different YOLO-TLA models with their YOLOv5 equivalents.
4.7 Comparison with state-of-the-arts
We also compared the performance of the two YOLO-TLA models with other state-of-the-art object detection models. To ensure a comprehensive and fair comparison, we divided the comparisons into two parts based on the number of parameters: YOLO-TLAs versus other lightweight models and YOLO-TLAm versus other large-scale models. The experimental results are presented in Tables 7 and 8, respectively.
The mAP@0.5:0.95 of YOLO-TLAs increased by 0.3% compared to YOLOx-s, with a reduction in computation by 1.5 GFLOPs, though the number of parameters increased by 0.49M. Compared to YOLOv5s, the mAP@0.5:0.95 of YOLO-TLAs improved by 4%, with an increase of only 2.26M in parameters. The mAP@0.5:0.95 of YOLO-TLAs is 7.5% higher than that of EfficientDet-D1. Although EfficientDet-D1 has a lightweight architecture, the reduction in model complexity has impacted accuracy.
YOLO-TLAm also performs better than YOLOx-m of the same scale, with a mAP@0.5:0.95 increase of 0.4% and a decrease in computation by 0.7 GFLOPs. Additionally, compared to the DETR model, although the mAP@0.5:0.95 of YOLO-TLAm only improved by 4.7%, its computation and parameter count are significantly lower than DETR.
5 Conclusion
This study tackles prevalent challenges in object detection communities and introduces a novel method, YOLO-TLA. This method demonstrates superior detection performance compared to the baseline YOLOv5, particularly in accurately identifying small objects. YOLO-TLA enhances YOLOv5 by integrating a tiny object detection layer into its neck network, and by incorporating a global attention mechanism to increase accuracy. To balance enhanced detection with model efficiency, a lightweight strategy is employed, incorporating the C3CrossCovn module in the backbone network to decrease model complexity. This strategy not only lowers complexity but also boosts accuracy, as confirmed by experimental results. Furthermore, we also successfully apply these improvements to the larger YOLOv5m model, creating the YOLO-TLAm model, which outperforms YOLOv5m in accuracy and stability. The results validate that the proposed enhancements are effective for larger models as well.